OpenStudy (anonymous):

Cos(x)/1+tan(x) - sin(x)/1+cot(x) = cos(x)-sin(x)

OpenStudy (sohailiftikhar):

?

OpenStudy (anonymous):

Try using the identities\[\tan x = \frac{ \sin x }{ \cos x }\] and \[\cot x = \frac{ \cos x }{ \sin x }\]